# Computing Lower and Upper Hitting Probabilities for Imprecise Markov Chains

### v1

@@ 4.2 Computing Upper Hitting Probabilities
$=p_{n}(x_{k^{*}}),$
where we used $\sum_{z\in\mathcal{X}}T_{n+1}(x_{k^{*}},z)=1$ as $T_{n+1}$ is a transition matrix. Since $T_{n}p_{n}](x_{k^{*}})=p_{n}(x_{k^{*}})$, we find $T_{n+1}p_{n}](x_{k^{*}})\leq[T_{n}p_{n}](x_{k^{*}})$ which forms a contradiction with our choice of $k^{*}$. Hence, there exists at least one $z\in\mathcal{S}$ such that $T_{n+1}(x_{k^{*}},z)>0$. Moreover, since $x\to^{T_{n+1}}x_{k^{*}}$ and $T_{n+1}(x_{k^{*}},z)>0$, we have $x\to^{T_{n+1}}z$.
So far, we have established that $x\to^{T_{n+1}}z$, with $p_{n}(x)<p_{n}(z)$. If $z\in A$, then $x\notin\mathcal{C}_{T_{n+1}}$ and we are done. Otherwise, we are back in the situation where we started, as we now want to establish that $z\to^{T_{n+1}}A$. The difference is that we have a strictly higher ‘current hitting probability’ of the state we are considering, as $p_{n}(x)<p_{n}(z)$. The state $z$ is again in one of the two cases we have already described: if it is in the first case, then $z\to^{T_{n+1}}y$ for some $y\in A$ and therefore also $x\to^{T_{n+1}}y$, which means that we can again conclude that $x\notin\mathcal{C}_{T_{n+1}}$. Otherwise, we find some $w\in\mathcal{X}$ such that $p_{n}(x)<p_{n}(z)<p_{n}(w)$ and $z\to^{T_{n+1}}w$. Since for every ‘iteration’ we do in this way we either conclude that $x\notin\mathcal{C}_{T_{n+1}}$ or strictly increase the hitting probability of the state under consideration, we must eventually be able to conclude that $x\notin\mathcal{C}_{T_{n+1}}$. Indeed, we cannot remain forever in the second case, since there are only finitely many states and each inequality in our sequence of considered hitting probabilities $p_{n}(x)<p_{n}(z)<p_{n}(w)<\dots$ is strict, meaning that we cannot repeat states. Hence we can finally conclude that $x\notin\mathcal{C}_{T_{n+1}}$, finishing the proof. ∎
@@
Let us first establish that the sequence $(p_{n})_{n\in\mathbb{N}}$ is well-defined. Since each $p_{n}$ is a vector of hitting probabilities, we have that $p_{n}\in[0,1]^{{\mathcal{X}}}$ for all $n\in\mathbb{N}$. Moreover, for any $n\in\mathbb{N}$, the minimal nonnegative solution in $0,1]^{{\mathcal{X}}}$ to
$p_{n}=\mathbb{1}_{A}+\mathbb{1}_{A^{c}}\cdot T_{n}p_{n}$
exists due to Proposition 2.5. Furthermore, for any $n\in\mathbb{N}$, the existence of the extreme point $T_{n+1}$ is guaranteed by Proposition 2.18.
We move on to showing that $(p_{n})_{n\in\mathbb{N}}$ is convergent and nondecreasing. We know that $p_{n}(x)=1$ for $x\in A$ for all $n\in\mathbb{N}$. Therefore $(p_{n})_{n\in\mathbb{N}}$ is trivially nondecreasing and convergent on $A$. The same holds for $\mathscr{W}_{\mathcal{T}}$}: for all $x\in{\mathscr{W}_{\mathcal{T}}}$ and for all $n\in\mathbb{N}$ it holds that $p_{n}(x)=0$ as $p_{n}$ is the hitting probability associated with $T_{n}$ and $A$ cannot be reached with any $T\in\mathcal{T}$ if we start from $x\in{\mathscr{W}_{\mathcal{T}}}$. This means that $(p_{n})_{n\in\mathbb{N}}$ is trivially nondecreasing and convergent on $\mathscr{W}_{\mathcal{T}}$}.
We continue by showing that $(p_{n})_{n\in\mathbb{N}}$ is nondecreasing on $A^{c}\setminus{\mathscr{W}_{\mathcal{T}}}$. Consider any $n\in\mathbb{N}$. Note that $T_{n+1}p_{n}=\overline{T}p_{n}\geq T_{n}p_{n}$ and that Lemma 4.4 yields $\mathscr{W}_{\mathcal{T}}=\mathcal{C}_{T_{n}}$ for all $n\in\mathbb{N}$. Then we have that
@@
where we used Lemma 2.6 for the last equation. Repeatedly expanding this inequality $m\in\mathbb{N}$ times yields
@@
Taking $m\to+\infty$ and applying Lemma 2.12 yields
@@
where the last step follows from the fact that $p_{n+1}=\mathbb{1}_{A}+\mathbb{1}_{A^{c}}\cdot T_{n+1}p_{n+1}$. Thus, we conclude that the sequence $(p_{n})_{n\in\mathbb{N}}$ is nondecreasing.
Since $(p_{n})_{n\in\mathbb{N}}$ is nondecreasing and $p_{n}\in[0,1]^{{\mathcal{X}}}$ for all $n\in\mathbb{N}$, it follows that the limit $p^{*}\coloneqq\lim_{n\to+\infty}p_{n}$ exists and $p^{*}\in[0,1]^{{\mathcal{X}}}$. We continue by showing that $p^{*}$ satisfies the system in Equation (23). Fix any $n\in\mathbb{N}$. Then
@@
Equation (24) now yields that
@@
where we used Lemma 2.16(3) for the second and third inequalities. Since all summands on the final right-hand side of the equation vanish as $n$ goes to infinity, we find that $p^{*}$ solves the system in Equation (23).
Recall that $p_{n}\in[0,1]^{{\mathcal{X}}}$, $p_{n}(x)=1$ for $x\in A$ and $p_{n}(x)=0$ for $x\in\mathscr{W}_{\mathcal{T}}$ for all $n\in\mathbb{N}$. Hence also $p^{*}\in[0,1]^{{\mathcal{X}}}$, $p^{*}(x)=1$ for all $x\in A$ and $p^{*}(x)=0$ for all $x\in\mathscr{W}_{\mathcal{T}}$. Due to Proposition 3.12, we can then conclude that $p^{*}$ is the unique minimal solution to the system in Equation (19) and so $p^{*}=\overline{p}^{\mathcal{T}}$.
@@ 4.2.1 Remark - the Tiebreaking Condition
Proposition 4.3 is very similar to Proposition 4.1, but imposes an additional condition on the next transition matrix:
@@
Initialize the algorithm at the extreme point
@@
which has $\mathcal{C}_{T_{1}}=\mathscr{W}_{\mathcal{T}}=\{1\}$. This matrix maximises the hitting probability having $p_{1}=\left(0,1,1\right)^{\top}$. Any transition matrix imposing $q_{1}=0$ achieves the maximum in $\max_{T\in\mathcal{T}}[Tp_{1}](2)=q_{2}+q_{3}=1$, so one possible choice for the extreme point $T_{2}$ is
@@ 5 Practical Considerations and Complexity Analysis
It remains to discuss how the algorithms described in Proposition 4.1 and Proposition 4.3 can be used in practice. This consists of three main aspects, namely how many iterations are required to reach convergence, the complexity of any such iteration, and how to find suitable starting points $T_{1}\in\mathcal{T}$. We assume that the set $\mathcal{T}$ is explicitly given, while the operators $\underline{T}$ and $\overline{T}$ are implicitly defined through Definition 2.17. This is the situation most common in practice, as the set $\mathcal{T}$ often characterises the uncertainty surrounding the true process by providing error bounds on the true transition probabilities. We mainly follow Krak’s complexity discussion [krak2021comphit].
@@ 5.1 Time per Iteration
The computation performed for each iteration in Proposition 4.1 and Proposition 4.3 can be divided into two parts. Firstly, we need to find the next element in the sequences $\left(p_{n}\right)_{n\in\mathbb{N}}$. By Proposition 2.13, this requires us to solve a linear system of equations, which can be done in at most $\mathcal{O}\left(N^{c}\right)$ time [Demmel2007]: the time required to multiply two $N\times N$ matrices. Currently, the best (published) estimate of $c$ is $c=2.371339$ [alman2025more].
Secondly, the next element of the sequences $\left(T_{n}\right)_{n\in\mathbb{N}}$ in these propositions needs to be computed. Hence, we investigate the complexity of finding some extreme point $T\in\mathcal{T}$ such that $\underline{T}f=Tf$ for some arbitrary $f\in\mathbb{R}^{\mathcal{X}}$, or, alternatively, finding some extreme point $T^{\prime}\in\mathcal{T}$ such that $\overline{T}f=T^{\prime}f$. The complexity of finding a suitable extreme point requires solving the optimisation problem from Definition 2.17 and therefore strongly depends on the definition of $\mathcal{T}$. Typically, finding such an extreme point will take at least $\mathcal{O}(N^{3})$ time [krak2021comphit]. In the remainder of the analysis, we will denote the complexity of finding an extreme point $T\in\mathcal{T}$ such that $\underline{T}f=Tf$ or $\overline{T}f=Tf$ for some arbitrary $f\in\mathbb{R}^{\mathcal{X}}$ by $\mathcal{O}(\operatorname{Extr}(\mathcal{T},\mathcal{X}))$. For both Proposition 4.1 and Proposition 4.3, the time per iteration is then $\operatorname{most}\mathcal{O}\left(N^{c}+\operatorname{Extr}(\mathcal{T},\mathcal{X})\right)$.
@@ 5.2 Starting Points
Let us discuss how to find a suitable $T_{1}\in\mathcal{T}$ for Proposition 4.1 and Proposition 4.3.
@@ 5.2.1 Lower Hitting Probabilities
Having obtained $\mathcal{A}_{\mathcal{T}}$, it remains to find a matrix $T_{1}\in\mathcal{T}$ such that $\mathcal{C}_{T_{1}}=\mathcal{A}_{\mathcal{T}}$. We observe that the matrix $T\in\mathcal{T}$ found in the last iteration of the algorithm satisfies $T\mathbb{1}_{\mathcal{D}_{k^{*}}}](x)=0$ for all $x\in\mathcal{A}_{\mathcal{T}}$. In summary, identifying $\mathcal{A}_{\mathcal{T}}$ and then finding $T_{1}\in\mathcal{T}$ to use in Proposition 4.1 requires at most $\mathcal{O}\left(\operatorname{Extr}(\mathcal{T},\mathcal{X})N\right)$ time.
@@ 5.2.2 Upper Hitting Probabilities
Let us now discuss how to find a suitable $T_{1}\in\mathcal{T}$ for Proposition 4.3. The matrix $T_{1}\in\mathcal{T}$ with $\mathcal{C}_{T_{1}}=\mathscr{W}_{\mathcal{T}}$ can be constructed row-wise. In particular, for every $x\in\mathcal{X}$, choose any $T_{1}(x,\cdot)$ in the relative interior of $\mathcal{T}_{x}$, taken with respect to the affine hull of $\mathcal{T}_{x}$. In this way, $T_{1}(x,y)>0$ for all $y\in\mathcal{X}$ such that $\overline{T}\mathbb{1}_{\{y\}}](x)>0$. The computational cost of this operation is $\mathcal{O}(N)$.
@@ 5.3 Number of Iterations
Let us now continue by investigating the number of iterations needed for convergence. We start by defining a sequence of vectors $\left(p_{n}\right)_{n\in\mathbb{N}}$ to be strictly decreasing if $p_{n+1}\leq p_{n}$ and $p_{n+1}\neq p_{n}$ for all $n\in\mathbb{N}$ and strictly increasing if $p_{n+1}\geq p_{n}$ and $p_{n+1}\neq p_{n}$. We then have the following result.
@@
Let $(p_{n})_{n\in\mathbb{N}}\subset[0,1]^{\mathcal{X}}$ be the sequence constructed in Proposition 4.1.Then $(p_{n})$ is either strictly decreasing at every coordinate, or there exists $n^{*}\in\mathbb{N}$ such that $p_{n+1}<p_{n}$ for all $n<n^{*}$ and $p_{k}=p_{*}$ for all $k\geq n^{*}$, where $p_{*}=\lim_{n\to\infty}p_{n}$.
Similarly, let $(p_{n})_{n\in\mathbb{N}}\subset[0,1]^{\mathcal{X}}$ be the sequence constructed in Proposition 4.3. Then $(p_{n})$ is either strictly increasing at every coordinate, or there exists $n^{*}\in\mathbb{N}$ such that $p_{n+1}>p_{n}$ for all $n<n^{*}$ and $p_{k}=p_{*}$ for all $k\geq n^{*}$, where $p_{*}=\lim_{n\to\infty}p_{n}$.
@@
Completely analogous to the proof of Krak’s Corollary 6 [krak2021comphit]. ∎
@@
Let $\mathcal{T}$ have finitely many extreme points; write $m\in\mathbb{N}$ for their number. Then the sequence $(p_{n})_{n\in\mathbb{N}}$ constructed in Proposition 4.1 stabilises in at most $m$ steps: there exist $n\in\mathbb{N}$ with $n\leq m$ such that $p_{k}=p_{*}$ for all $k\geq n$.
@@
Completely analogous to the proof of Krak’s Corollary 7 [krak2021comphit]. ∎
@@
Figure 1: Credal set of the first row of $\mathcal{T}(6)$.
Given $k\in\mathbb{N}$ and $n=n(k)=2^{k}$, we find the vector of hitting probabilities associated with $T_{n(k)}$ equal to $p_{n(k)}=(\frac{n}{n+1},1,0)$. We argue that the matrix $T_{n(k+1)}\in\mathcal{T}(m)$ satisfies $T_{n(k+1)}=\operatorname*{argmax}_{T\in\mathcal{T}}Tp_{n(k)}$, i.e.
@@
Initialize the algorithm for upper hitting probabilities with the matrix $S_{1}\coloneqq T_{n(1)}$. After one iteration, we get $S_{2}=T_{n(2)}$, as $S_{2}p_{n(1)}=\overline{T}p_{n(1)}$. Thus, only after cycling over all $m$ extreme points of $\mathcal{T}(m)$ we eventually find the maximising matrix in $S_{m}=T_{n(m)}$.

### v2

@@ 4.2 Computing Upper Hitting Probabilities
$=p_{n}(x_{k^{*}})\underset{=1}{\cancel{\sum_{z\notin\mathcal{S}}T_{n+1}(x_{k^{*}},z)}},$
where we used $p_{n}(z)\leq p_{n}(x_{k^{*}})$ for all $z\notin\mathcal{S}$. Since $T_{n}p_{n}](x_{k^{*}})=p_{n}(x_{k^{*}})$, we find $T_{n+1}p_{n}](x_{k^{*}})\leq[T_{n}p_{n}](x_{k^{*}})$ which is a contradiction. Therefore, there exists at least one $z\in\mathcal{S}$ such that $T_{n+1}(x_{k^{*}},z)>0$. Moreover, since $x\to^{T_{n+1}}x_{k^{*}}$ and $T_{n+1}(x_{k^{*}},z)>0$, we have $x\to^{T_{n+1}}z$, for some $z\in\mathcal{S}$.
So far, we have established that $x\to^{T_{n+1}}z$, with $p_{n}(x)<p_{n}(z)$. If $z\in A$, then $x\notin\mathcal{C}_{T_{n+1}}$ and the proof is complete. Otherwise, we find ourselves again in the starting situation, as we want to show that $z\to^{T_{n+1}}A$. The difference is that we have a strictly higher hitting probability of the state we are considering. The state $z$ is again in one of the two cases we have already described: if it is in the first case, then $z\to^{T_{n+1}}y$ for some $y\in A$ and therefore also $x\to^{T_{n+1}}y$, which implies $x\notin\mathcal{C}_{T_{n+1}}$. Otherwise, we find some $w\in\mathcal{X}$ with $w\neq x$ such that $p_{n}(x)<p_{n}(z)<p_{n}(w)$ and $z\to^{T_{n+1}}w$. Since we either establish $x\notin\mathcal{C}_{T_{n+1}}$ or we strictly increase the hitting probability, using the finiteness of the state space $\mathcal{X}$, we are able to conclude that $x\to^{T_{n+1}}A$ and $x\notin\mathcal{C}_{T_{n+1}}$. ∎
@@
The sequence $(p_{n})_{n\in\mathbb{N}}$ is well-defined since $p_{n}=\mathbb{1}_{A}+\mathbb{1}_{A^{c}\setminus\mathcal{C}_{T_{n}}}\cdot T_{n}p_{n}$ has a unique solution (Proposition 2.13) and the existence of the extreme point $T_{n+1}$ is guaranteed by Proposition 2.16.
By Proposition 2.13, we have that, for all $n\in\mathbb{N}$, $p_{n}(x)=1$ for all $x\in A$ and $p_{n}(x)=0$ for all $x\in\mathcal{C}_{T_{n}}=\mathscr{W}_{\mathcal{T}}$, therefore $(p_{n})_{n\in\mathbb{N}}$ is trivially convergent on $A\cup\mathscr{W}_{\mathcal{T}}$.
Next, we show that $(p_{n})_{n\in\mathbb{N}}$ is nondecreasing and convergent on $A^{c}\setminus{\mathscr{W}_{\mathcal{T}}}$. Recall that $T_{n+1}p_{n}=\overline{T}p_{n}\geq T_{n}p_{n}$ for all $n\in\mathbb{N}$. This yields
@@
where we used Lemma 2.6 and the fact that $\mathscr{W}_{\mathcal{T}}=\mathcal{C}_{T_{n}}$ for all $n\in\mathbb{N}$. Repeatedly expanding this inequality $m\in\mathbb{N}$ times yield
@@
Letting $m\to+\infty$ and applying Lemma 2.12 yields
@@
where the last step follows from Proposition 2.13. Thus, the sequence $(p_{n})_{n\in\mathbb{N}}$ is nondecreasing.
Since $(p_{n})_{n\in\mathbb{N}}$ is nondecreasing and $p_{n}\in[0,1]^{{\mathcal{X}}}$ for all $n\in\mathbb{N}$, it follows that the limit $p^{*}\coloneqq\lim_{n\to+\infty}p_{n}$ exists.
We move on to showing that $p^{*}$ satisfies the system in Equation (23). Fix any $n\in\mathbb{N}$. Then
@@
[…]
for any vector norm $||\cdot||$. Continuing (24)
@@
[…]
where we used Lemma 2.17(3) for the second and third inequalities. Since all summands on the final right-hand side of (25) vanish as $n\to+\infty$, we find that $p^{*}$ is a solution of the system in Equation (23).
It remains to establish that $p^{*}=\overline{p}^{\mathcal{T}}$, i.e. that $p^{*}$ is the minimal nonnegative solution of the system in Equation (23). We already know that they agree on $A\cup{\mathscr{W}_{\mathcal{T}}}$, so fix an arbitrary state $x\in A^{c}\setminus{\mathscr{W}_{\mathcal{T}}}$. We also know that $\overline{p}^{\mathcal{T}}(x)=\sup_{T\in\mathcal{T}}p^{T}(x)$ thus $p_{n}(x)\leq\overline{p}^{\mathcal{T}}(x)$ as $p_{n}$ is the hitting probability associated with $T_{n}\in\mathcal{T}$. Since we have established that $p_{n}(x)\to p^{*}(x)$ we conclude that $p^{*}(x)\leq\overline{p}^{\mathcal{T}}(x)$, and, by minimality of $\overline{p}^{\mathcal{T}}$, $p^{*}=\overline{p}^{\mathcal{T}}$.
@@ 4.2.1 Remark - the Tiebreaking Condition
Proposition 4.3 is similar to Proposition 4.1, but imposes an additional condition on the next transition matrix:
@@
Initialize the algorithm with
@@
which has $\mathcal{C}_{T_{1}}=\mathscr{W}_{\mathcal{T}}=\{1\}$. This matrix maximises the hitting probability having $\overline{p}^{\mathcal{T}}=p_{1}=\left(0,1,1\right)^{\top}$. Any transition matrix imposing $q_{1}=0$ achieves the maximum in $\max_{T\in\mathcal{T}}[Tp_{1}](2)=q_{2}+q_{3}=1$, so one possible choice for the extreme point $T_{2}$ is
@@
4.3 Interpretation as an MDP and Connection with Policy Iteration
The proposed algorithms can be seen as a variation of the policy iteration algorithm for Markov Decision Processes (MDPs) [23, 25]. In particular, for every state $x\in\mathcal{X}$ the set $\mathcal{T}_{x}$ can be seen as a set of actions. Then, every row of a transition matrix $T$, $T(x,\cdot)$, is an admissible action in state $x$. Moreover, every transition matrix $T\in\mathcal{T}$ can be seen as a policy $\pi_{T}:\mathcal{X}\to\mathcal{T}$. Taking the reward function $r$ as the hitting indicator $\mathbb{1}_{A}$, we find that, for every policy $\pi_{T}$, the value function is the hitting probability $p^{T}$. The optimal value $v^{*}=\underline{p}^{\mathcal{T}}=\inf_{T\in\mathcal{T}}p^{T}$ satisfies the (undiscounted) Bellman optimality equation [25] (which is equivalent to Equation (15)). The optimal value and the optimal policy can be computed using a policy iteration algorithm [15] which, similar to the algorithm in Proposition 4.1, alternates between policy evaluation, i.e. solving the linear system, and policy improvement.
These connections between imprecise Markov chains and MDPs have also previously been described in the literature; see, for instance, comments by Krak et al. [19, 20] about these relations, as well as for interpretations of similar algorithms for IMCs as being versions of value- and policy iteration methods for MDPs.
@@ 5 Practical Considerations and Complexity Analysis
It remains to discuss how the algorithms described in Proposition 4.1 and Proposition 4.3 can be used in practice. This consists of three main aspects, namely how many iterations are required to reach convergence, the complexity of any such iteration, and the computational cost of finding suitable starting points $T_{1}\in\mathcal{T}$. We assume that the set $\mathcal{T}$ is explicitly given, while the operators $\underline{T}$ and $\overline{T}$ are implicitly defined through Definition 2.15. This is the situation most common in practice, as the set $\mathcal{T}$ often characterises the uncertainty surrounding the true process by providing error bounds on the true transition probabilities.
@@ 5.1 Time per Iteration
The computation performed for each iteration in Proposition 4.1 and Proposition 4.3 can be divided into two parts. Firstly, we need to find the next element in the sequences $\left(p_{n}\right)_{n\in\mathbb{N}}$. By Proposition 2.13, this requires us to solve a linear system of equations, which can be done in at most $\mathcal{O}\left(N^{c}\right)$ time [6]: the time required to multiply two $N\times N$ matrices. Currently, the best (published) estimate of $c$ is $c=2.371339$ [1].
Secondly, we investigate the complexity of computing the next element of the sequences $\left(T_{n}\right)_{n\in\mathbb{N}}$. Finding such an extreme point requires solving the optimisation problem from Definition 2.15 and, therefore, strongly depends on the definition of $\mathcal{T}$. Typically, this computation takes at least $\mathcal{O}(N^{3})$ time [20]. In the remainder of the analysis, we denote the complexity of finding an extreme point $T\in\mathcal{T}$ such that $\underline{T}f=Tf$ or $\overline{T}f=Tf$ for some arbitrary $f\in\mathbb{R}^{\mathcal{X}}$ by $\mathcal{O}(\operatorname{Extr}(\mathcal{T},\mathcal{X}))$. Then, for Proposition 4.1 and Proposition 4.3, the time per iteration is at most $\mathcal{O}\left(N^{c}+\operatorname{Extr}(\mathcal{T},\mathcal{X})\right)$.
@@ 5.2 Starting Points
We now discuss the complexity of finding a suitable $T_{1}\in\mathcal{T}$ for Proposition 4.1 and Proposition 4.3.
@@ 5.2.1 Lower Hitting Probabilities
Having obtained $\mathcal{A}_{\mathcal{T}}$, it remains to identify a matrix $T_{1}\in\mathcal{T}$ such that $\mathcal{C}_{T_{1}}=\mathcal{A}_{\mathcal{T}}$. In the last iteration of the algorithm, we find $\underline{T}\mathbb{1}_{\mathcal{D}_{k^{*}}}](z)=0$ for all $z\in\mathcal{A}_{\mathcal{T}}$. Hence, we also have $|\mathcal{A}_{\mathcal{T}}|$ transition matrices, one for each $z\in\mathcal{A}_{\mathcal{T}}$, satisfying $S_{z}\mathbb{1}_{\mathcal{D}_{k^{*}}}](z)=[\underline{T}\mathbb{1}_{\mathcal{D}_{k^{*}}}](z)=0$. We then construct $T_{1}$ row-wise by setting $T_{1}(z,\cdot)=S_{z}(z,\cdot)$. In summary, identifying $\mathcal{A}_{\mathcal{T}}$ and finding $T_{1}\in\mathcal{T}$ to use in Proposition 4.1 requires at most $\mathcal{O}\left(\operatorname{Extr}(\mathcal{T},\mathcal{X})N\right)$ time.
@@ 5.2.2 Upper Hitting Probabilities
The matrix $T_{1}\in\mathcal{T}$ with $\mathcal{C}_{T_{1}}=\mathscr{W}_{\mathcal{T}}$ can be constructed row-wise. In particular, for every $x\in\mathcal{X}$, choose any $T_{1}(x,\cdot)$ in the relative interior of $\mathcal{T}_{x}$, taken with respect to the affine hull of $\mathcal{T}_{x}$. In this way, $T_{1}(x,y)>0$ for all $y\in\mathcal{X}$ such that $\overline{T}\mathbb{1}_{\{y\}}](x)>0$. The computational cost of this operation is $\mathcal{O}(N)$.
@@ 5.3 Number of Iterations
We now investigate the number of iterations needed for convergence. We start by defining a sequence of vectors $\left(p_{n}\right)_{n\in\mathbb{N}}$ to be strictly decreasing if $p_{n+1}\leq p_{n}$ and $p_{n+1}\neq p_{n}$ for all $n\in\mathbb{N}$ and strictly increasing if $p_{n+1}\geq p_{n}$ and $p_{n+1}\neq p_{n}$. We then have the following result.
@@
Let $(p_{n})_{n\in\mathbb{N}}\subset[0,1]^{\mathcal{X}}$ be the sequence constructed in Proposition 4.1. Then $(p_{n})$ is either strictly decreasing everywhere, or there exists $n^{*}\in\mathbb{N}$ such that $(p_{n})$ is strictly decreasing for all $n<n^{*}$ and $p_{k}=p^{*}$ for all $k\geq n^{*}$, where $p^{*}=\lim_{n\to\infty}p_{n}$.
Similarly, let $(p_{n})_{n\in\mathbb{N}}\subset[0,1]^{\mathcal{X}}$ be the sequence constructed in Proposition 4.3. Then $(p_{n})$ is either strictly increasing everywhere, or there exists $n^{*}\in\mathbb{N}$ such that $(p_{n})$ is strictly increasing for all $n<n^{*}$ and $p_{k}=p^{*}$ for all $k\geq n^{*}$, where $p^{*}=\lim_{n\to\infty}p_{n}$.
@@
Completely analogous to the proof of Krak’s Corollary 6 [20]. ∎
@@
Let $\mathcal{T}$ have $m\in\mathbb{N}$ extreme points. Then the sequence $(p_{n})_{n\in\mathbb{N}}$ constructed in Proposition 4.1 stabilises in at most $m$ steps: there exist $n^{*}\in\mathbb{N}$ with $n^{*}\leq m$ such that $p_{k}=p^{*}$ for all $k\geq n^{*}$.
@@
Completely analogous to the proof of Krak’s Corollary 7 [20]. ∎
@@
Figure 1: Credal set of $\mathcal{T}_{1}(6)$.
Given $k\in\mathbb{N}$ and $n=n(k)=2^{k}$, we find that the vector of hitting probabilities associated with $T_{n(k)}$ equals $p_{n(k)}=(\nicefrac{{n}}{{n+1}},1,0)$. We claim that the matrix $T_{n(k+1)}$ satisfies $T_{n(k+1)}=\operatorname*{argmax}_{T\in\mathcal{T}}Tp_{n(k)}$ for all $1\leq k\leq m-1$, i.e.
@@
Initialize the algorithm for upper hitting probabilities with the matrix $S_{1}\coloneqq T_{n(1)}$. After one iteration, we get $S_{2}=T_{n(2)}$, as
$S_{2}p_{n(1)}=\max_{T\in\mathcal{T}}Tp_{n(1)}=\overline{T}p_{n(1)},$
and so on. Thus, only after cycling over all $m$ extreme points of $\mathcal{T}(m)$ we eventually find the maximising matrix in $S_{m}=T_{n(m)}$.